A clown at a birthday party has brought along a helium cylinder, with which he intends to fill balloons. When full, each balloon contains 0.00210 m3 of helium at an absolute pressure of 1.20 x 10^5 Pa. The cylinder contains helium at an absolute pressure of 2.30 x 10^7 Pa and has a volume of 0.00360 m3. The temperature of the helium in the tank and in the balloons is the same and remains constant. What is the maximum number of people who will get a balloon?
1
Expert's answer
2020-09-25T09:11:27-0400

Use Boyle's law since temperature is kept constant which is given by


"P_1V_1=P_2V_2"


"1.20\times 10^{5}\times0.00210\times b=2.30\times 10^{7}\times0.00360"


where b is the number of balloons


"b=\frac{2.30\times10^{7}\times0.00360}{1.20\times10^{5}\times 0.00210}=328.57"


Maximum number of people who will get a balloon= 329